Draw and name all of the structural isomers that are ketones with five carbon atoms in 26. its longest chain and the moleculaI need help with the answers but also with the explanation and how to get there
Answer #1

26) Degree of unsaturation = C-H/2+1 = 5-10/2 +1 = 1

So, only one double bond is present and cyclic ring is not present.

Possible ketones are :

ir cs2-pentanone CamScanner 3-pentanone

Yes, aldehyde is also possible with this formula. Structure of aldehyde is :

No pentanal

Yes, ether is also possible. But that also contain atleast a double bond because degree of unsaturation is one.

27) Pent-2-ene on hydration give pentan-2-ol.

H, OH فیلم H0 115 111111 OH یک

28) 1-propanol has high boiling point due to strong H-BONDING whereas propanone has dipole-dipole interaction which is weaker than H-BONDING.